OpenStudy (love_to_love_you):

Help! Will fan and medal! What is a polynomial function in standard form with zeroes 1, 2, 3, and -3? a. x^4-3x^3+7x^2-27x-18 b.x^4+3x^3-7x^2-27x-18 c. x^4-3x^3-7x^2+27x-18 d. x^4+3x^3-7x^2-27x-18

OpenStudy (love_to_love_you):

@campbell_st @satellite73

OpenStudy (anonymous):

Sorry their both offline and I dont know this

OpenStudy (anonymous):

they're*

OpenStudy (campbell_st):

Use the factor theorem which requires you so substitute the zero, x = 1 etc, to show f(1) = 0 if successful with the 1st substitution repeat with f(2), then f(3) and f(-3) so as an example using the last equation \[f(1) = 1^3 +3\times 1^3 - 7 \times 1^2 - 27 \times 1 - 18 = -48\] so that eliminates the last equation. the alternative is to expand is x = 1 is a zero, then (x - 1) is a factor (x -1)(x-2)(x-3)(x+3) hope it helps.. it is a little tedious